# Article Category

### v2

@@ 7.6 PROOF OF LEMMA 1: JACOBIAN OF BOUNDARY PROJECTION
$2\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}+\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}$ $=2I_{d}$ (51)
[…]
We solve this system by substitution starting from Equation˜54
$\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}=G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (55)
which we substitute into Equation˜51
$\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}=2I_{d}-2G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,.$ (56)
We substitute this into Equation˜52
$\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}=2G_{\mathcal{A}_{s}}^{T}-2G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (57)
with which we decouple Equation˜53
$(\lambda_{D}^{*}-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}\,.$ (58)
To solve this linear system, we utilise the KKT complementarity slackness conditions in element-wise notation
$\lambda_{i}^{*}(K_{i,:}z_{*}-h_{i})=\lambda_{i}^{*}(\gamma_{i}^{*}-1)=0\quad\forall i=1,\ldots,n\,.$ (59)
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=1$ to fulfil the conditions. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=1\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We examine the rows of Equation˜58
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ (60)
$=-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T})_{i,:}\,,$
where we utilise the diagonalised form of the optimal variables and denote full rows or columns with a colon subscript. Next, we combine this with the index sets
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}$ =0=0= 0 $i\in\mathcal{I}_{a}$ (61)
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=(G_{\mathcal{A}_{s}}^{T})_{i,:}$ $i\in\mathcal{I}_{i}\,.$ (62)
We partition the Jacobian according to the index sets $\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$. The assumption of strong duality implies strict complementarity $\lambda_{i}^{*}>0$. Therefore, Equation˜61 specifies the submatrix $\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}=\textbf{0}$. The left-hand side of Equation˜62 reduces to
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=\sum_{j}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$ (63)
$=\sum_{j\in\mathcal{I}_{i}}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$
$=\sum_{j\in\mathcal{I}_{i}}{\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{i,:}\left(G_{\mathcal{A}_{s}}\right)_{:,j}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}\,,$
using Equation˜61. We isolate the remaining submatrix
$\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}=(G_{\mathcal{A}_{s}}^{T})_{\mathcal{I}_{i},:}\,.$ (64)
The system is always solvable since $n\geq d$, and the right-hand side lies trivially in the column space of the left-hand side. We utilise the Moore-Penrose inverse to solve the system
[…]
$\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}$ $=\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\,.$ (66)
We plug this result into Equation˜55
@@
$\frac{da_{s}}{da}$ $=\begin{bmatrix}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{a}}&\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\end{bmatrix}\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$ (67)
@@
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}$
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}$
The resulting matrix is the projection matrix onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. It is also unique since any homogeneous part of the solution lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$, see Section˜7.7 for details.
The rank of a projection matrix is equal to the design matrix itself $\text{rank}(\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a})=\text{rank}(\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}})$. Due to the size of $G_{\mathcal{A}_{s}}$, the rank of the Jacobian is at most $d$ when all inequality constraints are inactive, which is only the case for interior points of the zonotope, i.e. safe actions. However, for boundary points, i.e. mapped unsafe actions, at least one constraint is active. Therefore, the rank is at most $d-1$.
Differentiability of Equation˜67 may not be given at points where the active constraint set $\mathcal{I}_{a}$ changes. However, these active set transitions occur on measure-zero sets, and the projection remains of Class $C^{0}$ everywhere. We employ any element from the Clarke subdifferentiable at non-differentiable points, ensuring our approach remains well-defined.
7.7 EXTENSION OF PROOF 1: JACOBIAN UNIQUENESS OF BOUNDARY PROJECTION
[…]
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}+\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}\right)\,.$
Using the property that $\text{Null}(A^{T}A)=\text{Null}(A)$, any homogeneous solution $\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}$ lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. Therefore, $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}=\textbf{0}$, and the expression simplifies to
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}\,,$
which depends only on the unique particular solution provided by the pseudoinverse.
7.8 PROOF OF LEMMA 2: CONVEXITY OF INDUCED SAFE ACTION SETS
@@
We start by inserting Equation˜10 into Equation˜2
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\in\mathcal{A}:Ma_{i}+\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\subseteq\mathcal{S}_{s}\right\}\,.$ (68)
Since applying the matrix $M$ is a linear operation and convexity preserving, the safe action set is convex if and only if the set of all translations $\mathcal{T}=\left\{t:t+\mathcal{Z}_{const.}\subseteq\mathcal{S}_{s}\right\}$ is convex. The definition of convexity dictates
$\lambda x+(1-\lambda)y\in\mathcal{S}_{s}$ (69)
for any $x,y\in\mathcal{S}_{s}$ and $\lambda\in\left[0,1\right]$. It also applies to $\mathcal{T}$, since for any $t_{1},t_{2}\in\mathcal{T}$ setting $x=\mathcal{Z}_{const.}+t_{1}$ and $y=\mathcal{Z}_{const.}+t_{2}$ yields
$\lambda(\mathcal{Z}_{const.}+t_{1})+(1-\lambda)(\mathcal{Z}_{const.}+t_{2})$ (70)
$=\mathcal{Z}_{const.}+(\lambda t_{1}+(1-\lambda)t_{2})\in\mathcal{S}_{s}\,,$
which is the respective convexity condition of $\mathcal{T}$.
7.9 PROOF OF LEMMA 3: FULL RANK JACOBIAN OF RAY MASK
{IEEEproof}
Since a ray mask does not modify the action in the centre, the Jacobian for this case is trivially the identity and of full rank. The easiest application and differentiation of the other case is in spherical coordinates, centred at the safe centre. We transform the coordinates with
@@ 7.9 PROOF OF LEMMA 3: FULL RANK JACOBIAN OF RAY MASK
where we adopt the convention of the radius being the first coordinate. In these coordinates, the ray mask modifies a single coordinate
$a_{s,o}=\begin{bmatrix}{\color[rgb]{0,0,0}\omega\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{a}}}a_{o,r}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}={\color[rgb]{0,0,0}\begin{bmatrix}\omega\lambda_{\mathcal{A}_{s}}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}}\,$ (72)
since $\lambda_{a}=a_{o,r}$ in this coordinate system. Finally, we transform the safe action back
@@
where the inverse function theorem [49] relates the Jacobians of the transformations as
$\frac{\partial a_{s}}{\partial a_{s,o}}=\frac{\partial a_{o}}{\partial a}^{-1}\,.$ (75)
This relation characterises a similarity transformation, which means $\frac{\partial a_{s}}{\partial a}$ and $\frac{\partial a_{s,o}}{\partial a_{o}}$ are similar and share the same eigenvalues. The Jacobian $\frac{\partial a_{s,o}}{\partial a_{o}}$ is
@@
$\sigma_{i}=\begin{cases}{\color[rgb]{0,0,0}\frac{\partial\omega}{\partial a_{o,r}}\lambda_{\mathcal{A}_{s}}}&\text{if }i=1\\ 1&\text{else}\,.\end{cases}$ (77)
Since they are all non-zero and the Jacobian is a square matrix, it has full rank.

### v3

@@ B PROOF OF THEOREM 1: JACOBIAN OF BOUNDARY PROJECTION
$2\frac{\partial a_{s}}{\partial a}+\frac{\partial\nu}{\partial a}$ $=2I_{d}$ (55)
${\begin{bmatrix}I_{n}&-I_{n}\end{bmatrix}}\frac{\partial\lambda}{\partial a}-G_{\mathcal{A}_{s}}^{T}\frac{\partial\nu}{\partial a}$ $=\textbf{0}$ (56)
$\lambda_{D}^{*}\begin{bmatrix}I_{n}\\ -I_{n}\end{bmatrix}\frac{\partial\gamma}{\partial a}+\begin{bmatrix}\gamma_{D}^{*}-I_{n}&\textbf{0}\\ \textbf{0}&-\gamma_{D}^{*}-I_{n}\end{bmatrix}\frac{\partial\lambda}{\partial a}$ $=\textbf{0}$ (57)
$\frac{\partial a_{s}}{\partial a}-G_{\mathcal{A}_{s}}\frac{\partial\gamma}{\partial a}$ $=\textbf{0}\,.$ (58)
We solve this system by substitution, starting from Equation˜58:
$\frac{\partial a_{s}}{\partial a}=G_{\mathcal{A}_{s}}\frac{\partial\gamma}{\partial a}\,$, (59)
which we substitute into Equation˜55
$\frac{\partial\nu}{\partial a}=2I_{d}-2G_{\mathcal{A}_{s}}\frac{\partial\gamma}{\partial a}\,.$ (60)
We substitute this into Equation˜56
$\frac{1}{2}{\begin{bmatrix}I_{n}&-I_{n}\end{bmatrix}}\frac{\partial\lambda}{\partial a}=G_{\mathcal{A}_{s}}^{T}-G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}}\frac{\partial\gamma}{\partial a}\,.$ (61)
To utilise the remaining Equation˜57, we partition it by the activity of the inequality constraint in Equation˜45c utilising the KKT complementarity slackness conditions [80, Eq. 4.3] in element-wise notation
$\forall i=1,\ldots,n:\lambda_{i}^{*}(K_{i,:}z_{*}-h_{i})=0\,.$ (62)
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=\pm 1$, where we write $\pm$ for shortness to denote the lower and upper part of the supremum norm. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=\pm\textbf{1}\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We reorder the generators according to the index sets and partition Equation˜57 into
@@
$\begin{bmatrix}\left(\pm\lambda^{*}_{\mathcal{I}_{a}}\right)_{D}\left(\frac{\partial\gamma}{\partial a}\right)_{\mathcal{I}_{a},:}\\ (\pm\gamma^{*}_{\mathcal{I}_{i}}-\textbf{1})_{D}\left(\frac{\partial\lambda}{\partial a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}=\begin{bmatrix}\textbf{0}\\ \textbf{0}\end{bmatrix}\,$, (63)
where we utilise the diagonalised form of the optimal variables and denote full rows or columns with a colon subscript. Using the assumption of strict complementarity we have
$\forall i\in\mathcal{I}_{a}:\lambda^{*}_{i}>0$ (64)
$\forall i\in\mathcal{I}_{i}:\pm\gamma^{*}_{i}-1\neq 0\,$, (65)
which means the diagonal matrices $\left(\lambda^{*}_{\mathcal{I}_{a}}\right)_{D}$ and $(\pm\gamma^{*}_{\mathcal{I}_{i}}-\textbf{1})_{D}$ are both invertible, yielding
$\left(\frac{\partial\gamma}{\partial a}\right)_{\mathcal{I}_{a},:}$ $=\textbf{0}$ (66)
$\left(\frac{\partial\lambda}{\partial a}\right)_{\mathcal{I}_{i},:}$ $=\textbf{0}\,.$ (67)
To obtain the remaining part $\left(\frac{\partial\gamma}{\partial a}\right)_{\mathcal{I}_{i},:}$, we examine the inactive rows in Equation˜61 and substitute Equation˜67 yielding
$\left(G^{T}_{\mathcal{A}_{s}}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{\partial\gamma}{\partial a}\right)_{\mathcal{I}_{i},:}=\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\,$, (68)
which is always solvable, since the right-hand side lies trivially in the column space of the left-hand side. We utilise the Moore-Penrose inverse as a solution
$\left(\frac{\partial\gamma}{\partial a}\right)_{\mathcal{I}_{i,:}}=\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}$ (69)
and combine Equation˜66 and Equation˜69 into
@@
$\frac{\partial\gamma}{\partial a}=\begin{bmatrix}\textbf{0}\\ \left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\end{bmatrix}\,$. (70)
Next, we insert Equation˜70 into Equation˜59
$\frac{\partial a_{s}}{\partial a}=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\,.$ (71)
The Jacobian is the orthogonal projection onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. Despite the non-uniqueness of the solution in Equation˜69, the resulting Jacobian in Equation˜71 is unique [81, Theorem 1], as the orthogonal projector on a subspace is unique. The rank of an orthogonal projection or projection matrix $P=X\left(X^{T}X\right)^{\dagger}X^{T}$ is equal to the design matrix $X$ itself $\text{rank}(\frac{\partial a_{s}}{\partial a})=\text{rank}(\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}})$ [82, Chapter 9]. Due to the shape of $G_{\mathcal{A}_{s}}\in\mathbb{R}^{d\times n}$, the rank of the Jacobian is at most $d$. However, the columns of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$ cannot span all of $\mathbb{R}^{d}$, since at the optimum the residual $a-a_{s}$ must be orthogonal to this span. Therefore, the rank is at most d1d-1.
C PROOF OF THEOREM 3: FULL RANK JACOBIAN OF RAY MASK
@@
An easy application and differentiation of the ray mask for $\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}>\epsilon$ is in spherical coordinates, centred at the safe centre. We transform the coordinates with
@@
where we adopt the convention of the radius being the first coordinate. In this coordinate system $c_{\mathcal{A}_{s}}=0$, $\lambda_{a}=a_{o,r}$, and $d_{a}$ is the first canonical unit vector, which means Equation˜21 modifies only the first coordinate
$a_{s,o}=\begin{bmatrix}\omega(a_{o,r},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})\lambda_{\mathcal{A}_{s}}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}\,$. (73)
Finally, we transform the safe action back
@@
where the inverse function theorem [49, Theorem 3.3.2] relates the Jacobians of the transformations as
$\frac{\partial a_{s}}{\partial a_{s,o}}=\left(\frac{\partial a_{o}}{\partial a}\right)^{-1}\,.$ (76)
This relation characterises a similarity transformation, which means $\frac{\partial a_{s}}{\partial a}$ and $\frac{\partial a_{s,o}}{\partial a_{o}}$ share the same eigenvalues. The Jacobian $\frac{\partial a_{s,o}}{\partial a_{o}}$ is
@@
$\sigma_{i}=\begin{cases}\frac{\partial\omega(a_{o,r},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})}{\partial a_{o,r}}\lambda_{\mathcal{A}_{s}}&\text{if }i=1\\ 1&\text{else}\,.\end{cases}$ (78)
Since they are all non-zero, recall the condition on a valid mapping $\frac{\partial w(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})}{\partial\lambda_{a}}>0$, and the Jacobian is a square matrix, it has full rank.
D LEARNING CURVES OF ALL LEARNING ALGORITHMS IN NON-SAFEGUARDED TRAINING
Figure˜7 shows the performance of SHAC, SAC, and PPO in the absence of any safeguarding. These results complement the main text by illustrating how each algorithm behaves under unconstrained conditions.
E LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH BOUNDARY PROJECTION
Figure˜8 provides the learning curves for SHAC, SAC, and PPO when applying the boundary projection method. This visualization highlights the effect of the projection on training stability and convergence.
F LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH RAY MASK
Figure˜9 presents the learning curves obtained with the ray mask. It illustrates the comparative performance of SHAC, SAC, and PPO when safeguarded by this technique.
Figure 9: Learning curves SHAC, SAC, and PPO with ray mask.
G LEARNING CURVES OF BOTH APPROXIMATIONS
Figure˜10 compares the safeguarded training results of SHAC under two different safe centre approximation methods. This allows us to assess the relative effectiveness of the ORP and ZRP approximations.